Two cells of emf E1 and E2 are shown,when a potentiometer is connected between A and B the balancing length of the potentiometer wire is 300cm.on connecting same potentiometer between A and C,the balancing length is 100cm.the ratio E1÷E2 is ?

Grade:12

2 Answers

Vikas TU
6 years ago
Dear Student,
​The question is here incomplete.
Pls do check the qstn. and Repost it again for the proper clarification.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)
Akash rawat
6 years ago
See guys according to question E1=k×300For second part it depends on the connection of cell if negative terminal on same side then E1+E2=k×100If in opposite direction thenE1-E2=k×100Hope u got it Akash RawatClass 12(jms)Dehradun